A 25kg block is pushed a distance 5m along a level floor at constant speed by a force at an angle of 30° below the horizontal. The coefficient of friction between the block and floor is 0.4 how much work is done?
1
Expert's answer
2017-01-26T13:18:12-0500
A=FScos30; Fcos30=kN=k (mg-Fsin30)=0.4 (250-0.5F)=100-0.2F; F (cos30+0.2)=100; F~100; A=100×5×0.86=430J